# Med-R3: Enhancing Medical Retrieval-Augmented Reasoning of LLMs via Progressive Reinforcement Learning

### v3

@@ Abstract
In medical scenarios, effectively retrieving external knowledge and leveraging it for rigorous logical reasoning is of significant importance. Despite their potential, existing work has predominantly focused on enhancing either retrieval or reasoning capabilities of the models in isolation, with little attention given to their joint optimization, which leads to limited coordination between the two processes. Additionally, current methods rely heavily on supervised fine-tuning (SFT), which can cause models to memorize existing problem-solving pathways, thereby restricting their generalization ability when confronted with novel problem contexts. Furthermore, while some studies have explored to improve retrieval-augmented reasoning in general domains via reinforcement learning, their reward function designs do not adequately capture the specific demands of the medical domain. To address these challenges, we introduce Med-R3, a Medical Retrieval-augmented Reasoning framework driven by progressive Reinforcement learning. In this framework, we first develop the model’s ability to perform logical reasoning over medical problems. Subsequently, on the basis of this foundation, we adaptively optimize the retrieval capability to better align with the characteristics of knowledge corpus and external information utilization throughout the reasoning process. Finally, we conduct joint optimization of the model’s retrieval and reasoning coordination. Extensive experiments indicate that Med-R3 could achieve state-of-the-art performances, with LLaMA3.1-8B-Instruct + Med-R3 surpassing closed-sourced GPT-4o-mini by 3.93% at a comparable parameter scale, while Qwen2.5-14B augmented with Med-R3 shows a more substantial gain of 13.53%.
Med-R3: Enhancing Medical Retrieval-Augmented Reasoning of LLMs via Progressive Reinforcement Learning
Keer Lu, Zheng Liang§, Youquan Li, Jiejun Tan§, Da Pan§, Shusen Zhang§, Guosheng Dong§, Zhonghai Wu, Bin Cui, Huang Leng, Wentao Zhang Center for Data Science, Academy for Advanced Interdisciplinary Studies, Peking University, School of CS & Key Lab of High Confidence Software Technologies, Peking University, §Baichuan Inc. {keer.lu, youquan.li}@stu.pku.edu.cn, {bin.cui, lenghuang, wentao.zhang}@pku.edu.cn
1 Introduction
With the rapid development of artificial intelligence, large language models (LLMs) have shown remarkable potential in various fields Lewkowycz et al. (2022). However, when applied to the medical domain, LLMs face unique challenges. Accurate diagnosis of medical conditions requires rigorous logical reasoning Lucas et al. (2024); Savage et al. (2024), yet unlike domains such as mathematics or programming, where internal knowledge is often sufficient, the complexity and specificity of medical diagnosis necessitate the integration of external, up-to-date, and domain-specific knowledge Xiong et al. (2024a); Lu et al. (2025). Therefore, during the process of solving medical problems, both retrieval and reasoning play crucial roles.
@@ 1 Introduction
Figure 1: Comparison of Med-R3 (bottom) with existing methods (top) for medical problem-solving.
@@
Despite their potential, existing works face some limitations, as depicted in Figure˜1: (C1) Limited Coordination between Retrieval and Reasoning: While research has been devoted to independently improving the retrieval Jeong et al. (2024); Xiong et al. (2024a, b) and reasoning Goh et al. (2024); Lucas et al. (2024) capabilities of models in the medical domain, they do not adequately align the retrieval strategy with requirements of the reasoning process, which is a critical gap that limits the system’s end-to-end performance Yoran et al. (2023); Asai et al. (2024). (C2) Lack of Generalization in SFT: Subsequent studies have investigated to achieve the joint optimization of these components during the supervised fine-tuning (SFT) stage Wang et al. (2024). However, recent findings suggest that SFT inherently causes models to memorize task-specific shortcuts rather than learning generalizable reasoning in novel scenarios Lai et al. (2025). Conversely, reinforcement learning (RL) has shown high effectiveness at enhancing sophisticated reasoning capabilities in LLMs Jaech et al. (2024); Guo et al. (2025); Team et al. (2025). (C3) Tailored Reward Design for Medical Scenarios: While recent works such as R1-Searcher Song et al. (2025) and ReSearch Chen et al. (2025) have explored training models for retrieval-augmented reasoning via reinforcement learning in general-domain tasks, their reward strategies are not well suited to the medical domain, where factors including the coverage of entities and relationships, as well as the credibility of retrieved documents, are of paramount importance in the reasoning process.
@@
To address these challenges, we introduce Med-R3, a Medical Retrieval-augmented Reasoning framework driven by progressive Reinforcement learning. For C1 and C2, we perform a progressive RL to co-optimize the model’s retrieval and reasoning capabilities in three distinct stages: (1) Stage 1: Reasoner Cultivation. We begin by developing the model’s logical reasoning abilities when addressing medical questions. (2) Stage 2: Retriever Awakening. Building upon the reasoning capabilities acquired in Stage 1, we adaptively optimize the retriever to better align with the retrieval features of the knowledge base and external information utilization during the model’s reasoning process. (3) Stage 3: Joint Optimization. Finally, we refine the coordination between retrieval and reasoning in models to enhance their collaborative performance in medical scenarios. For C3, we design specialized rewards tailored to the characteristics of medical reasoning at each training stage, encompassing aspects such as the coverage of entities and relations during the reasoning process, as well as the effectiveness of retrieval mechanisms that jointly consider the quality of medical evidence Sackett et al. (1996) and the influence of retrieved documents within the overall reasoning trajectory. The main contributions can be summarized as follows:
@@
  • Training Framework Advancement. We propose Med-R3, a novel training framework to improve the retrieval-augmented reasoning performance within the medical scenario based on progressive reinforcement learning, which jointly enhances the model’s capability in retrieval and reasoning through a three-stage co-optimization strategy.
@@
  • Performance and Effectiveness. Experiments indicate that models trained with Med-R3 significantly improve medical performances. Notably, LLaMA3.1-8B-Instruct + Med-R3 surpasses the closed-sourced proprietary model GPT-4o-mini by 3.93% at a comparable parameter scale, while Qwen2.5-14B integrated with Med-R3 shows a more substantial gain of 13.53%.
2 Med-R3
@@
Figure 2: The progressive reinforcement learning pipeline of Med-R3. We designed tailored rewards for different training stages to facilitate the effective optimization of the model’s ability to interpret and answer medical questions.
2.1 Training Data Construction
@@
Inspired by the data selection and verifiability transformation process of HuatuoGPT-o1 Chen et al. (2024b), we filtered medical questions based on their cognitive complexity and compatibility with reinforcement learning, and then constructed reference reasoning trajectories for verifying.
2.1.1 Data Selection
@@
We utilize both closed-set exam questions and rare disease diagnostic datasets for our progressive reinforcement learning (RL) phase, including the training sets from MedQA-USMLE, MedQA-MCMLE Jin et al. (2020), MedMCQA Pal et al. (2022), and RareArena [^1] For the RareArena dataset, we split the data into training and test sets with a ratio of 8:2 through random sampling. THUMedInfo (2025). Data filtering and processing are conducted referring to the following criteria:
@@
  • Content Suitability Judgment: We employed DeepSeek-V3 Liu et al. (2024) to exclude the medical questions that are unsuitable for RL or lacking unique answers (e.g., those requiring the identification of incorrect options).
@@
  • Reasoning Complexity Filtering: We then used pass@n metrics to assess the question difficulty. Specifically, we employed DeepSeek-R1 Guo et al. (2025) to generate up to $n=16$ answers along with the corresponding reasoning process for each medical question. DeepSeek-V3 was then utilized to evaluate both the intermediate reasoning steps and final answers, with scores ranging from 1 to 5. We filtered out questions that were either too simple (achieving full correctness within a few rollouts, $n^{\prime}\leq 3$) or excessively ambiguous (consistently receiving scores of 2 or lower across all 16 rollouts). Finally, we obtained a training dataset comprising 2,140 questions from MedQA-USMLE, 1,204 from MedQA-MCMLE, 6,748 from MedMCQA, and 429 from the RareArena dataset.
2.1.2 Construction of Reference Reasoning Trajectories
@@
We performed data augmentation by rephrasing closed-set questions into open-ended formats, generating high-quality reasoning process, and further extracting structured reasoning paths that capture the pivotal logical steps.
@@
  • Open-Ended Standardization: We then reformatted the multiple-choice questions into open-ended formats using DeepSeek-V3 Liu et al. (2024), transforming questions with predefined answer choices into tasks that require free-form, reasoning-based responses.
@@
  • Reasoning Process Generation: Similar to the rollout process of data selection based on reasoning complexity, we prompted DeepSeek-R1 to generate all possible reasoning paths along with final answers for each medical question. During the reasoning process, the model dynamically retrieves relevant external knowledge as needed and integrates the retrieved information to support continued inference. Then we employed DeepSeek-V3 as an evaluator to assess the generated reasoning processes and answers on a scale from 1 to 5, retaining those with a score of 5. This procedure yielded 5 high-quality reasoning processes per question on average.
@@
    $\mathcal{G}=\{(h,l,t,s)\mid h\in\mathcal{E},l\in\mathcal{L},t\in\mathcal{E},s\in\mathbb{B}\}$ (1)
    where $\mathcal{E}$ denotes the set of all entities, $\mathcal{L}$ represents the set of all relations, and $(h,l,t)$ is a triple in which the head entity $h$ is connected to the tail entity $t$ via the relation $l$. For each triplet, we add a binary attribute $s$, where $\mathbb{B}=\{0,1\}$, specifying whether it was obtained through external knowledge retrieval ($s=1$) or internally generated via the model’s reasoning capabilities ($s=0$).
2.2 Progressive Reinforcement Learning
We employ a three-stage pipeline to enhance the model’s capabilities in retrieval and reasoning for medical problems, as illustrated in Figure˜2.
2.2.1 Stage 1: Cultivating the Reasoner
In this stage, we focus on developing the model’s ability to reason through medical problems. Unlike previous works Chen et al. (2024b); Liu et al. (2025a) that typically verify model performance based only on the correctness of the final answer, we incorporate a reasoning process evaluation mechanism into the reward function, aligning closely with the characteristics of clinical decision-making. Overall, the reward function design in Stage 1 is the sum of normalized components:
@@ 2.2.1 Stage 1: Cultivating the Reasoner
Reasoning Process.  This aspect constitutes a core component in fostering the model’s deliberative reasoning capabilities during Stage 1. Given the characteristics of medical reasoning, which requires a holistic understanding of interconnected entities and their relationships—such as those between symptoms, diagnoses, and therapeutic interventions Wu et al. (2023); Gao et al. (2025), we design the following reward for evaluation:
$\mathcal{R}_{reasoning}=\tilde{R}_{semantic}+\tilde{R}_{statistic}+\tilde{R}_{logical}$ (4)

### v4

@@ Abstract.
In medical scenarios, effectively retrieving external knowledge and leveraging it for rigorous reasoning is important. Despite their potential, existing work has predominantly focused on enhancing either retrieval or reasoning capabilities in isolation, with little attention given to their joint optimization, which leads to limited coordination between the two processes. Additionally, current methods rely heavily on supervised fine-tuning (SFT), which can cause models to memorize existing problem-solving pathways, thereby restricting their generalization ability when confronted with novel problem contexts. Furthermore, while some studies have explored to improve retrieval-augmented reasoning in general domains via reinforcement learning, their reward function designs do not adequately capture the specific demands of the medical domain. To address these challenges, we introduce Med-R3, a Medical Retrieval-augmented Reasoning framework driven by progressive Reinforcement learning. In this framework, we first develop the model’s ability to perform logical reasoning over medical problems. Subsequently, based on this foundation, we adaptively optimize the retrieval capability to better align with the characteristics of knowledge corpus and external information utilization throughout the reasoning process. Finally, we conduct joint optimization of the model’s retrieval and reasoning coordination. Experiments indicate that Med-R3 could achieve state-of-the-art performances, with LLaMA3.1-8B-Instruct + Med-R3 surpassing closed-sourced GPT-4o-mini by 3.93% at a comparable parameter scale, while Qwen2.5-14B augmented with Med-R3 shows a more substantial gain of 13.53%.
Retrieval-Augmented Reasoning, Reinforcement Learning, Knowledge Graph Construction, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@
Figure 1. Comparison of Med-R3 (bottom) with existing methods (top) for medical problem-solving.
1. Introduction
With the rapid development of artificial intelligence, large language models (LLMs) have shown remarkable potential in various fields (Lewkowycz et al., 2022). However, when applied to the medical domain, LLMs face unique challenges. Accurate diagnosis of medical conditions requires rigorous logical reasoning (Lucas et al., 2024; Savage et al., 2024), yet unlike domains such as mathematics or programming, where internal knowledge is often sufficient, the complexity and specificity of medical diagnosis necessitate the integration of external, up-to-date, and domain-specific knowledge (Xiong et al., 2024a; Lu et al., 2025). Therefore, during the process of solving medical problems, both retrieval and reasoning play crucial roles.
Figure 2. The pipeline of training data construction for Med-R3.
@@ 1. Introduction
Despite their potential, existing works face some limitations, as depicted in Figure 1: (C1) Limited Coordination between Retrieval and Reasoning: While research has been devoted to independently improving the retrieval (Jeong et al., 2024; Xiong et al., 2024a, b) and reasoning (Goh et al., 2024; Lucas et al., 2024) capabilities of models in the medical domain, they do not adequately align the retrieval strategy with requirements of the reasoning process, which is a critical gap that limits the system’s end-to-end performance (Yoran et al., 2023; Asai et al., 2024). (C2) Lack of Generalization in SFT: Subsequent studies have investigated to achieve the joint optimization of these components during the supervised fine-tuning (SFT) stage (Wang et al., 2024). However, recent findings suggest that SFT inherently causes models to memorize task-specific shortcuts rather than learning generalizable reasoning in novel scenarios (Lai et al., 2025). Conversely, reinforcement learning (RL) has shown high effectiveness at enhancing sophisticated reasoning capabilities in LLMs (Jaech et al., 2024; Guo et al., 2025; Team et al., 2025). (C3) Tailored Reward Design for Medical Scenarios: While recent works such as R1-Searcher (Song et al., 2025) and ReSearch (Chen et al., 2025b) have explored training models for retrieval-augmented reasoning via reinforcement learning in general-domain tasks, their reward strategies are not well suited to the medical domain, where factors including the coverage of entities and relationships, as well as the credibility of retrieved documents, are of paramount importance in the reasoning process.
@@
To address these challenges, we introduce Med-R3, a Medical Retrieval-augmented Reasoning framework driven by progressive Reinforcement learning. For C1 and C2, we first construct a medical training set of approximately 10K instances, each containing reference reasoning trajectories, and then we perform a progressive RL to co-optimize the model’s retrieval and reasoning capabilities in three distinct stages: (1) Stage 1: Reasoner Cultivation. We begin by developing the model’s logical reasoning abilities when addressing medical questions. (2) Stage 2: Retriever Awakening. Building upon the reasoning capabilities acquired in Stage 1, we adaptively optimize the retriever to better align with the retrieval features of the knowledge base and external information utilization during the model’s reasoning process. (3) Stage 3: Joint Optimization. Finally, we refine the coordination between retrieval and reasoning in models to enhance their collaborative performance in medical scenarios. For C3, we design specialized rewards tailored to the characteristics of medical reasoning at each training stage, encompassing aspects such as the coverage of entities and relations during the reasoning process, as well as the effectiveness of retrieval mechanisms that jointly consider the quality of medical evidence (Sackett et al., 1996) and the influence of retrieved documents within the overall reasoning trajectory. The main contributions are three-fold:
@@
  • Training Framework Advancement. We propose Med-R3, a training framework to improve the retrieval-augmented reasoning performance in the medical field based on progressive reinforcement learning, which jointly enhances the model’s capability in retrieval and reasoning via a three-stage optimization strategy.
@@
  • Performance and Effectiveness. Extensive experiments indicate that models trained with Med-R3 significantly improve medical performances. Notably, LLaMA3.1-8B-Instruct + Med-R3 surpasses the closed-sourced proprietary model GPT-4o-mini by 3.93% at a comparable parameter scale, while Qwen2.5-14B integrated with Med-R3 shows a more substantial gain of 13.53%.
2. Med-R3
@@
Figure 3. The progressive reinforcement learning pipeline of Med-R3. We designed tailored rewards for different training stages to facilitate the effective optimization of the model’s ability to interpret and answer medical questions.
2.1. Training Data Construction
@@
Inspired by the data selection and verifiability transformation process of HuatuoGPT-o1 (Chen et al., 2025a), we filtered medical questions based on their cognitive complexity and compatibility with the feature of reinforcement learning, and then constructed reference reasoning trajectories for verification, as illustrated in Figure 2.
2.1.1. Data Selection
@@
We utilize both closed-set exam questions and rare disease diagnostic datasets for our reinforcement learning phase, including the training sets from MedQA-USMLE, MedQA-MCMLE (Jin et al., 2020), MedMCQA (Pal et al., 2022), and RareArena [^1] For the RareArena dataset, we split the data into training and test sets with a ratio of 8:2 through random sampling. (THUMedInfo, 2025). Data filtering and processing are conducted referring to the following criteria:
@@
  • Content Suitability Judgment: We employed the frontier model, e.g., DeepSeek-V3 (Liu et al., 2024), to exclude the medical questions that are unsuitable for RL or lacking unique answers (e.g., those requiring the identification of incorrect options).
@@
  • Reasoning Complexity Filtering: We then used pass@n metrics to assess the question difficulty. Specifically, we employed DeepSeek-R1 (Guo et al., 2025) to generate up to $n=16$ answers along with the corresponding reasoning process for each medical question. DeepSeek-V3 was then utilized to evaluate both the intermediate reasoning steps and final answers, with scores ranging from 1 to 5. We filtered out questions that were either too simple (achieving full correctness within a few rollouts, $n^{\prime}\leq 3$) or excessively ambiguous (consistently receiving scores of 2 or lower across all 16 rollouts). Finally, we obtained a training dataset comprising 2,140 questions from MedQA-USMLE, 1,204 from MedQA-MCMLE, 6,748 from MedMCQA, and 429 from the RareArena dataset.
2.1.2. Construction of Reference Reasoning Trajectories
@@
We performed data augmentation by rephrasing closed-set questions into open-ended formats, generating high-quality reasoning process, and further extracting structured reasoning paths that capture the pivotal logical steps leading to the final answer.
@@
  • Open-Ended Standardization: We then reformatted the multiple-choice questions into open-ended formats using DeepSeek-V3 (Liu et al., 2024), transforming questions with predefined answer choices into tasks that require free-form, reasoning-based responses.
@@
  • Reasoning Process Generation: Similar to the rollout process of data selection based on reasoning complexity, we prompted DeepSeek-R1 to generate all possible reasoning paths along with final answers for each medical question. During the reasoning process, the model dynamically retrieves relevant external knowledge as needed and integrates the retrieved information to support continued inference. Then we employed DeepSeek-V3 as an evaluator to assess the generated reasoning processes and answers on a scale from 1 to 5, retaining those with a score of 5. Ultimately, for each medical question, we obtained an average of 5 reference reasoning processes.
@@
    (1) $\mathcal{G}=\{(h,l,t,s)\mid h\in\mathcal{E},l\in\mathcal{L},t\in\mathcal{E},s\in\mathbb{B}\}$
    where $\mathcal{E}$ denotes the set of all entities, $\mathcal{L}$ represents the set of all relations, and $(h,l,t)$ is a triple in which the head entity $h$ is connected to the tail entity $t$ via the relation $l$. For each triplet, we add a binary attribute $s$, where $\mathbb{B}=\{0,1\}$, specifying whether it was obtained through external retrieval ($s=1$) or internally generated via the model’s reasoning capabilities ($s=0$).
2.2. Progressive Reinforcement Learning
We employ a three-stage progressive pipeline to enhance the model’s capabilities in retrieval and reasoning for addressing problems in the medical scenario, as illustrated in Figure 3.
2.2.1. Stage 1: Cultivating the Reasoner
In this stage, we focus on developing the model’s ability to reason through medical problems. Unlike previous works (Chen et al., 2025a; Liu et al., 2025a) that typically verify model performance based only on the correctness of the final answer, we incorporate a reasoning process evaluation mechanism into the reward function, aligning closely with the characteristics of clinical decision-making. Overall, the reward function design in Stage 1 is the sum of normalized components:
@@ 2.2.1. Stage 1: Cultivating the Reasoner
Reasoning Process.  This aspect constitutes a core component in fostering the model’s reasoning capabilities during Stage 1. Given the characteristics of medical reasoning, which requires a holistic understanding of interconnected entities and their relationships, e.g., those between symptoms, diagnoses, and therapeutic interventions (Wu et al., 2023; Gao et al., 2025), we design the following reward for evaluation:
(4) $\mathcal{R}_{reasoning}=\tilde{R}_{semantic}+\tilde{R}_{statistic}+\tilde{R}_{logical}$